f $20,000 is invested in a savings account offering 3.5% per year, compounded continuously, how fast is the balance growing after 4 years? (Round your answer to the nearest cent.)
$ 1 per yr
Answer #1
The effective interest rate with continuous componding is
i = e^(.035)-1 = 3.562%

After four years the balance is
(with continuous compounding)
20,000*(1.0362)^4 = $23,005.48.

The balance is growing at an instantaneous rate of
0.035*23,005.48 = $805.19 /year

but over the year to follow, the earnings will be
$832.80 /yr
answered by: Jonathan
Answer #2
thanks
answered by: Florence